# Reinan Lima1, Graziella Bedenik2, Farshad Yazdani1, Matthew Robertson2, and Elyson Ádan Nunes Carvalho1 1Universidade Federal de Sergipe, São Cristóvão, Brazil

### v1

@@ I Introduction
The study and modelling of planar capacitors remain highly relevant due to their broad applications [1, 2, 3, 4, 5, 6]. They play a crucial role in microelectronics and sensing technologies, where their low cost, fast response, and flexible electrode design make them attractive for sensor development [7, 8]. Additionally, unintended parasitic capacitances — often formed by PCB traces behaving like planar capacitors — can significantly affect the performance of various electronic devices [9]. Therefore, understanding and accurately modelling these components is essential for intentional design and performance optimization.
However, accurate modelling of planar capacitors presents significant challenges due to intrinsic high-frequency effects, which include not only parasitic inductance but also ohmic losses in the copper and dielectric losses in the substrate [10, 11]. These phenomena render the conventional parallel-plate formula ($C=\epsilon_{0}\frac{A}{d}$) inadequate, since planar capacitors exhibit substantially different field distributions [12], as illustrated in Fig. 1. Consequently, two main approaches emerge in the literature: 1) employing empirical methods that adjust existing formulas through correction terms and constants [13, 14, 15], and 2) developing completely new formulations that explicitly incorporate high-frequency effects. Although more complex, this second approach provides greater accuracy and physical insight.
@@
This work proposes the development of a new simplified planar capacitor model consisting of two identical rectangular plates, which incorporates the effects of both the substrate and the medium between the conductor traces on the total capacitance. The primary objective is to develop a model that accurately represents the behaviour of these components at frequencies below 1 GHz, while incorporating the effects of parasitic inductances and losses associated with the copper and substrate. To validate the proposed model, a comparative study will be conducted between the theoretical results and Vendik et al. ’s model, as well as experimental data measured from real planar capacitors. This analysis enables the evaluation of the model’s accuracy and application limits and helps identify potential areas for future work.
The structure of this work is organized as follows: in section II, we present in detail the mathematical development of the proposed model, including all formulations and approximations considered. In section III, the experiments carried out to validate the proposed model are presented. Section IV discusses the obtained results, with comparative analyses between the model predictions and experimental data. Finally, section V draws the main conclusions about the effectiveness of the developed model and its limitations, and proposes potential model improvements and complementary investigations that could expand its applicability.
@@ II-A Capacitance
$\small C=C_{1}+C_{2}=\varepsilon_{0}l\ln\left(\dfrac{2W}{d}+1\right)\left[\frac{\varepsilon_{sub}}{\pi}+\frac{\varepsilon_{r}}{\pi-\varphi_{0}}\right],$ (4)
@@ III Experimental Methodology
The experiments aim to validate the proposed series RLC model presented in (12) and its corresponding capacitance formula, represented in (4) and (5), through systematic comparisons with the established model by Vendik et al. [19] and experimental measurements on real planar capacitors. The complete methodology is detailed in subsections III-A and III-B.
@@ IV Results and Discussion
The results illustrated in Fig. 5 show distinct behaviours between the two analytical models regarding substrate permittivity variation. The proposed model exhibits a noticeable shift in resonance frequency when replacing the FR-4 substrate ($\varepsilon_{r}=4.3$) with Teflon ($\varepsilon_{r}=2.1$). This displacement agrees with the expected reduction in effective capacitance, since capacitance is directly proportional to the dielectric medium’s permittivity. The proposed model’s sensitivity to material properties demonstrates its ability to accurately capture the substrate’s influence on capacitor behaviour, which is a key design aspect, as selecting substrates with different dielectric properties would yield significantly different responses if this parameter weren’t adequately considered. In contrast, Vendik’s model shows practically unchanged response to the same material variation, demonstrating its limitation in adequately incorporating substrate effects. This behavioural difference highlights the advantage of the proposed model, which more completely considers field lines traversing both the substrate and the space between copper traces, thereby capturing the actual influence of dielectric properties on component performance. The proposed model’s ability to appropriately respond to these variations demonstrates its superior performance and utility for designs where substrate material selection is a relevant design parameter.
@@ V Conclusion
By doing so, we establish an important theoretical milestone by proposing a more comprehensive formulation that integrates three fundamental aspects: enhanced capacitance calculation (including substrate and inter-conductor medium effects), parasitic inductance modelling, and resistive loss consideration. This unified approach serves as a foundation for future research and practical applications in high-frequency circuit design, providing valuable tools for the design and analysis of planar capacitors.

### v2

@@ I Introduction
The study and modeling of planar capacitors remain highly relevant due to their widespread use across various applications, particularly in microelectronics and sensing technologies [1, 2, 3, 4, 5, 6]. Their low cost, fast response, and flexible electrode configurations make them especially attractive for sensor development [7, 8]. At the same time, unintentional parasitic capacitances—often introduced by PCB traces that behave like planar capacitors—can significantly influence the performance of electronic circuits and measurement systems [9]. As a result, a deeper understanding and accurate modelling of planar capacitors is crucial not only for intentional sensor design but also for minimizing unwanted parasitic effects.
Despite being a well-established topic, planar capacitor modelling still faces significant challenges, particularly when accounting for measurement phenomena at frequencies in the range of tens to hundreds of MHz. Many existing models oversimplify the physical behaviour by neglecting subtle yet critical effects such as parasitic inductance, ohmic losses in conductive traces, and dielectric losses in the substrate material [10, 11]. These factors limit the applicability of the classical parallel-plate capacitor equation ($C=\epsilon_{0}\frac{A}{d}$), which assumes uniform field distributions that do not hold in planar geometries [12], as shown in Fig. 1. To address these limitations, three main strategies have emerged in the literature: (1) empirical corrections to traditional formulas through fitting parameters and adjustment factors [13, 14, 15]; (2) numerical simulations (e.g., finite element method or finite difference time domain approaches) that solve Maxwell’s equations directly for complex geometries, and (3) the development of new formulations that explicitly incorporate high-frequency and geometric effects. Although the latter is more complex, it offers improved accuracy and physical insight, making it particularly valuable for modern high-frequency applications. Numerical simulations, while accurate, often obscure fundamental parametric relationships within localized results, making it difficult to extract general trends—a limitation that analytical models overcome.
@@
This work introduces a simplified yet comprehensive analytical model for planar capacitors, developed to support the simulation and analysis of measurement systems operating below 1 GHz. The primary contribution of this paper is the development of an analytical model that incorporates the effects of both the substrate and the medium between conductor traces on the total capacitance of planar capacitors—factors often overlooked or treated separately in previous work. We consolidate and extend several equation fragments from the literature into a unified analytical expression capable of estimating the capacitor’s impedance. While numerical simulations can yield higher accuracy, the proposed analytical model offers key advantages: (i) it reveals explicit relationships between physical parameters, enabling identification of dominant effects across frequency ranges; (ii) it provides a computationally efficient starting point for early-stage design, particularly when multi-physics simulations are too resource-intensive, and (iii) it enables predicting the influence of geometry and dielectric properties on the component, which is especially beneficial for sensing applications.
@@ II-A Capacitance
$\small C=C_{1}+C_{2}=\varepsilon_{0}l\ln\left(\dfrac{2W}{d}+1\right)\left[\frac{\varepsilon_{sub}}{\pi}+\frac{\varepsilon_{r}}{\pi-\varphi}\right],$ (4)
@@ III Experimental Methodology
The experimental validation focuses on the proposed series RLC model defined in (12), along with its associated capacitance expressions (4) and (5). To assess the model’s accuracy, we conducted systematic comparisons against both the established model by Vendik et al. [19] and experimental measurements from real planar capacitors. As this study emphasizes analytical modelling, electromagnetic simulations were not included; instead, the validation relies directly on physical component measurements. The full methodology is described in subsections III-A and III-B.
@@ IV Results and Discussion
The results illustrated in Fig. 5 show distinct behaviours between the two analytical models regarding substrate permittivity variation. The proposed model exhibits a noticeable shift in resonance frequency when replacing the FR-4 substrate ($\varepsilon_{r}=4.3$) with Teflon ($\varepsilon_{r}=2.1$). This displacement agrees with the expected reduction in effective capacitance, since capacitance is directly proportional to the dielectric medium’s permittivity. The proposed model’s sensitivity to material properties demonstrates its ability to accurately capture the substrate’s influence on capacitor behaviour, which is a key design aspect, as selecting substrates with different dielectric properties would yield significantly different responses if this parameter weren’t adequately considered. In contrast, Vendik’s model shows practically unchanged response to the same material variation, demonstrating its limitation in adequately incorporating substrate effects. This behavioural difference highlights the advantage of the proposed model, which more completely considers field lines traversing both the substrate and the space between copper traces, thereby capturing the actual influence of dielectric properties on component performance. The proposed model’s ability to appropriately respond to these variations demonstrates its superior performance and utility for designs where substrate material selection is a relevant design parameter, which reinforces our initial argument that existing models often neglect subtle, but essential effects, explored in section I.
@@ V Conclusion
By doing so, we establish an important theoretical milestone by proposing a more comprehensive formulation that integrates three fundamental aspects: enhanced capacitance calculation (including substrate and inter-conductor medium effects), parasitic inductance modelling, and resistive loss consideration. This unified approach serves as a foundation for future research and practical applications in high-frequency circuit design, providing valuable tools for the design and analysis of planar capacitors and allowing the study of the phenomenon and its physical meaning directly from the modelled parameters.